OpenStudy (anonymous):

The velocity of sound on a particular day outside is 331 meters/second. What is the frequency of a tone if it has a wavelength of 0.6 meters?

OpenStudy (anonymous):

To determine the frequency of sound on a day of a given temperature, we need to use the following equation\[v = \lambda f\] where lambda is the wavelength of the wave and f is the frequency of the sound wave. Rearranging this equation algebraically for frequency yields\[f = \frac{ v }{ \lambda }\] To solve for the frequency we divide the velocity of sound by the wavelength or:\[f = \frac{ 331 m/s }{ 0.6 m }=551.7 s ^{-1} = 552 Hz\] As you can see, 1/s is defined as a Hz (Hertz).

OpenStudy (anonymous):

Thank you :)